Robert gave up his job earning $70,000/year and converts a house that he owns into an art museum. Before opening his museum, he had been renting the house out for $25,000/year. Robert hired two employees for his museum at a total cost of $50,000/yr, and had to pay $10,000/yr for utilities. The museum brought in total revenues of $100,000 the first year. Someone even offered him $900,000 for one of his paintings but he declined this offer.

1. What is Robert's economic profit? Explain your answer.
Answer #1

Economic profit = Total revenue - (explicit cost + implicit cost)

Explicit cost is the out of pocket cost of doing a business, while implicit cost is the opportunity cost of doing that business.

Robert's Total revenue = $100,000

Explicit cost = $(50,000 + 10,000) = $60,000

Robert's implicit cost or opportunity cost is the income he gave up to own his art museum and the rent he could have received if he didn't convert his house into a museum. Also, the $900,000 is his opportunity cost of keeping the paintings instead of selling it.

Implicit cost = $(900,000 + 70,000 + 25,000) = $995,000

Economic profit = $100,000 - ($60,000 + 995,000) = - $955,000